OpenStudy (anonymous):

Harvey plans to use 5 ft of shelving for four shelves whose lengths are to be a series of consecutive even numbers. How many inches long should each shelf be?

OpenStudy (raden):

1 ft = 12 inc, so 5 ft = 60 inc let that numbers are : x, x+2, x+4,x+6 x+x+2+x+4+x+6 = 60 4x+12=60 4x=60-12 4x=48 x=12 thus, each shelf be : 12,14,16,18 (in inc)